No book available as this is a home quiz. Have looked for help on the web to help my son but cannot find anything to help teach this concept. Haven't done alegebra in 30 years :) 4x+3y=12
Answer by ptaylor(1894)

No book available as this is a home quiz. Have looked for help on the web to help my son but cannot find anything to help teach this concept. Haven't done alegebra in 30 years :) 4x+3y=12

It's not exactly clear to me what the concept is that you are trying teach your son.
The equation that you give is the equation of a straight line. Now the standard form for this is the slope-intercept form:
y=mx+b where m is the slope of the line and b is the y-intercept.
We can put the equation into slope-intercept form as follows:
4x+3y=12 subtract 4x from both sides
4x-4x+3y=12-4x or
3y=-4x+12 divide through by 3
y=-(4/3)x+4
slope is -4/3 and the y intercept is 4

If this is not what you are looking for, I suggest that you re-post.

Hope this helps------ptaylor